# SGPMIL: Sparse Gaussian Process Multiple Instance Learning

### v1

@@ Abstract
Multiple Instance Learning (MIL) offers a natural solution for settings where only coarse, bag-level labels are available, without having access to instance-level annotations. This is usually the case in digital pathology, which consists of gigapixel sized images. While deterministic attention-based MIL approaches achieve strong bag-level performance, they often overlook the uncertainty inherent in instance relevance. In this paper, we address the lack of uncertainty quantification in instance-level attention scores by introducing SGPMIL, a new probabilistic attention-based MIL framework grounded in Sparse Gaussian Processes (SGP). By learning a posterior distribution over attention scores, SGPMIL enables principled uncertainty estimation, resulting in more reliable and calibrated instance relevance maps. Our approach not only preserves competitive bag-level performance but also significantly improves the quality and interpretability of instance-level predictions under uncertainty. SGPMIL extends prior work by introducing feature scaling in the SGP predictive mean function, leading to faster training, improved efficiency, and enhanced instance-level performance. Extensive experiments on multiple well-established digital pathology datasets highlight the effectiveness of our approach across both bag- and instance-level evaluations. Our code will be made publicly available.
@@ 1 Introduction
Building on recent progress in MIL for WSI analysis, embedding-based MIL models with attention pooling have demonstrated strong performance at the bag level while offering interpretable heatmaps that highlight task-relevant tissue regions [17, 34, 45]. These methods rely on attention mechanisms to weigh instance contributions, yet the resulting scores are often interpreted heuristically as indicators of instance relevance. In this work, we introduce a probabilistic formulation of the attention mechanism by learning a posterior distribution over attention weights. This formulation enables principled prediction uncertainty estimation and enhances the interpretability of instance-level predictions, while maintaining strong bag-level performance.
@@
Recent attention-based MIL aggregation schemes have demonstrated impressive bag-level performance across various metrics [17, 25, 34, 45], on a variety of datasets. These results have been further enhanced by leveraging recent foundation models, which provide powerful instance-level representations. The aforementioned MIL works showcase plotted attention heatmaps of test set cases for explainability purposes, in which the provided attention heatmaps seem to focus on instances of interest which, for example, in a binary classification setting indicate the presence or absence of a positive class instance. While the attained bag-level performance is impressive, attention values are often interpreted as class-specific instance discriminators. We hypothesize that explicitly modeling uncertainty over these attention scores—by learning a posterior distribution—can improve instance-level performance, while preserving strong bag-level accuracy.
@@
Probabilistic deep learning offers principled tools for uncertainty estimation and improved model calibration [12]. This is particularly valuable in MIL, where probabilistic frameworks may clarify both prediction confidence, statistical properties as well as the inner workings of pooling operations [42, 13]. Bayesian extensions of instance-based MIL have been explored to quantify uncertainty, particularly under data scarcity and label noise [13, 28]. Furthermore, bayesian extensions of the embedding-based MIL have explored the performance/explainability gains that one has by inducing uncertainty to the attention scoring mechanism and imposing prior regularization [42, 33]. These approaches typically incorporate variational inference and Monte Carlo (MC) sampling to model distributional uncertainty at the instance and bag level. Moreover, current challenges in probabilistic MIL approaches include numerical instability during training [33], lack extensive evaluation on multiple datasets and tasks with various encoders or require apriory decision on the range of instance spatial correlations in the scoring function [42].
@@
In this paper, we introduce SGPMIL, an efficient and robust probabilistic attention-based MIL framework grounded in Sparse Gaussian Processes. By designing a novel probabilistic formulation, our main contributions are structured along the following axes:
@@
(i) A fast and efficient probabilistic framework. SGPMIL builds upon prior work on Sparse Gaussian Processes for MIL and introduces three key innovations: (1) a feature scaling term in the predictive mean of the learnable attention SGP variational posterior, which improves representational flexibility and adaptation to varying input scales; (2) a more stable and robust sampling strategy from the variational posterior during training, addressing numerical instabilities in Gaussian Process-based MIL approaches; and (3) an interpretable instance-level attention normalization mechanism across attention samples, enabling more consistent and meaningful instance-level predictions under the MIL setting.
@@
Extensive experiments across multiple datasets demonstrates on-par bag-level performance on binary classifications tasks, with SGPMIL outperforming leading MIL approaches on multiclass datasets. Furthermore, we provide both intuitive motivation and empirical validation for the hypothesis that the inducing point approach in our MIL formulation captures salient histological morphologies and effectively focuses on task-relevant instances, leading to superior instance-level performance.
@@ 2.1 Deterministic frameworks
Many embedding-based MIL approaches compute instance-level attention scores through fixed parametric transformations. Ilse et al. [17] introduce attention-based deep MIL (ABMIL), where the slide-level representation is computed as a weighted average of instance embeddings. The attention weights are derived via a gated mechanism, which applies a two-layer neural network combining a $\tanh$ and sigmoid activation followed by a dot product, allowing the model to learn instance relevance in a differentiable, end-to-end manner. Lu et al. in Clustering-constrained attention MIL (CLAM) [25] extended the gated attention mechanism by introducing an extra clustering objective via a smooth SVM loss to constrain the patch feature space to being more separable with respect to class characterizing evidence, in various histopathology datasets while releasing a comprehensive toolbox for data pre-processing and model training. Li et al. in Dual-Stream MIL Network (DSMIL) [22] propose a dual-attention mechanism that combines instance-level and bag-level attention, where the most critical instance is selected via a classifier stream and its features guide the aggregation of the remaining patches. This is coupled with multiscale feature extraction and contrastive pretraining, leading to improved tumor detection performance. Shao et al. in Transformer-based MIL (TransMIL) [34] make use of two self-attention heads [39] to weigh the instances and introduce spatial correlations using a novel Pyramid Position Encoding Generator (PPEG) achieving strong performance in bag-level classification metrics. More recent work like DGRMIL [45] model instance correlation by introducing cross-attention across instances and a set of learnable tokens which are trained by a positive instance alignment mechanism and a novel diversity loss term. Despite strong bag-level performance, existing attention based MIL methods rarely assess how well attention mechanisms identify task-relevant instances. Being deterministic, they also lack the capacity to estimate predictive uncertainty—crucial in medical applications. We address these gaps through systematic instance-level evaluation and demonstrate improved instance identification.
@@ 2.2 Probabilistic frameworks
AGP [33] integrates a Sparse Gaussian Process (SGP) layer within a multi-layer perceptron (MLP) to model attention distributions over instances for classification tasks. Given input embeddings, the SGP layer computes a variational multivariate normal posterior distribution over the attention values associated with each instance in a bag; Monte Carlo (MC) sampling yields multiple attention values, which are used to reweigh the instance embeddings, resulting in multiple samples of bag-level representations that are subsequently classified. The model is trained by maximizing the Evidence Lower Bound (ELBO) with respect to the model parameters. Another related approach is BayesMIL [42], which addresses the interpretability of attention mechanism in MIL by modeling probabilistic instance weights. Specifically, they introduce multiplicative Gaussian noise [20], into the projection layers and the classifier. Furthermore, they propose the Slide-Dependent Patch Regularizer (SDPR), which imposes a prior encouraging attention scores to remain close to zero for negative bags, while allowing values near zero or one for positive bags in binary classification settings. To capture spatial correlations among instance attention distributions, they incorporate a conditional random field (CRF) on top of the learned attention distributions. Despite their promise, existing probabilistic approaches lack evaluations with pathology foundation encoders on cancer datasets, extensive instance-level evaluation on other datasets or rely on fixed spatial smoothing kernels that introduce task-specific inductive biases and require manual tuning [42]. AGP, in particular, faces numerical instabilities and scalability issues on larger datasets, and does not assess instance-level performance. In this work, we address these limitations by introducing a feature scaling term in the mean of the SGP variational posterior distribution, an efficient sampling strategy during training and a in AGP, addressing different instabilities, while conducting a comprehensive evaluation of both bag- and instance-level performance on the widely used public benchmarks.
@@
Gaussian Processes (GPs) offer a powerful flexible and elegant framework for regression, as is indicated by an extensive line of works [35, 19, 37]. As non-parametric standalone regressors, they mitigate the need for application-specific model architectures and a priori decision on model complexity [18, 41]. Furthermore, [30, 21] showcase GPs equivalence to infinite-width single-layer fully-connected neural network with an i.i.d. prior over its parameters and with infinitely wide deep networks respectively which in turn are known to be universal function approximators [16]. Yet, these benefits come at the cost of time and storage complexity of the order of $\mathcal{O}(n^{3})$ and $\mathcal{O}(n^{2})$ respectively mainly due to inverting a $n\times n$ kernel matrix, $n$ being the number of training points, which is prohibitive for large datasets [29]. To overcome the time and storage cost bottleneck, Sparse Gaussian Processes were originally introduced by [35] and subsequently extended to allow for mini-batch training [15]. We leverage the scalable variational GPs framework introduced in [15] to induce uncertainty over instance attention distributions, and incorporate distributional correlation through kernel-based terms, showing that modeling such semantic correlations substantially improves instance-level performance.
@@ 3.1 Preliminaries
Figure 1: SGPMIL architecture overview. On the left, we illustrate the complete processing pipeline: WSIs are segmented and tiled into patches, which are then encoded using a self-supervised pathology foundation model (UNI [5]). The resulting patch embeddings are passed through an attention-based MIL head incorporating a probabilistic SGP mechanism. The attention-weighted embeddings are summed and projected through a trainable MLP. On the right, we highlight the probabilistic attention component. The SGP layer receives the projected embeddings along with learnable inducing point locations and infers a variational posterior over patch-level attention scores. Multiple attention samples are drawn to reweigh the patch embeddings, which are then aggregated into stochastic slide-level representations and classified via a linear layer followed by softmax.
Multiple Instance Learning (MIL)
In MIL, the input is a bag of instances, $X_{i}=\{x_{1},\ldots,x_{K_{i}}\}$ with k-th instance $x_{k}\in\mathbb{R}^{D}$, associated with a bag label $Y_{i}\in\mathbb{N}$, while the instance labels $\{y_{1},\ldots,y_{K_{i}}\}$ remain unobserved during training; here $K_{i}$ denotes the number of instances in bag $i$. A dataset consists of $N$ bag-label pairs, denoted as $\mathcal{D}={(X_{i},Y_{i})}_{i=1}^{N}$, and let $\mathcal{C}=\{1,\ldots,C\}$ denote the set of class labels. The standard MIL assumption is defined as:
$Y_{i}=\begin{cases}0,&\text{if }\sum_{k=1}^{K_{i}}y_{k}=0\\[5.0pt] c,&\text{if }\exists\,i_{k}\text{ such that }y_{k}=c,c\in\mathcal{C}\end{cases}$ (1)
Sparse Gaussian Processes (SGPs)
Following standard notation, given a training dataset $\{\left(X_{i},Y_{i}\right)\}_{i=1}^{n}$ of size $n$, where $Y_{i}=F_{i}+\epsilon_{i}$ a noisy observation of the noise-free real-valued latent function $F(\cdot):X\in\mathcal{X}\rightarrow\mathbb{R}$ evaluated at $X_{i}$ and $\epsilon_{i}\sim\mathcal{N}\left(0,\sigma^{2}\right)$ models independent observation noise, an additional set of m learnable inducing points $\mathcal{Z}=\{Z_{i}\}_{i=1}^{m}$, $m\ll n$ are introduced to summarize the information of the entire dataset. These points $Z_{i}\in\mathcal{Z}$ reside in the same space as the inputs $X_{i}\in\mathcal{X}$ and are optimized during training to capture the underlying data distribution. In our formulation, n is the total number of instances across all bags, m is the number of learnable prototypical instance representations and we train the SGP layer using batches of instances, where each batch contains instances from the same bag. Let the vector $\mathbf{F}\in\mathbb{R}^{n}$ contain the values of function $F$ evaluated at each training point $X_{i}$ and the vector $\mathbf{U}\in\mathbb{R}^{m}$ be the values of the same function at the points $Z_{i}\in\mathcal{Z}$. Assuming that the vectors $(\mathbf{F},\mathbf{U})$ are jointly gaussian, the SGP model evaluated at finite sets of inputs $(\mathbf{X},\mathbf{Z})$ can be characterized by the following equations,
@@
$p(\mathbf{F},\mathbf{U})$ $=\mathcal{N}\left(\begin{bmatrix}\mathbf{F}\\ \mathbf{U}\end{bmatrix}\middle|\begin{bmatrix}\boldsymbol{\mu}_{\mathbf{X}}\\ \boldsymbol{\mu}_{\mathbf{Z}}\end{bmatrix},\begin{bmatrix}\mathbf{\Sigma}_{XX}&\mathbf{\Sigma}_{XZ}\\ \mathbf{\Sigma}_{ZX}&\mathbf{\Sigma}_{ZZ}\end{bmatrix}\right)$ (3)

### v2

@@ Abstract
Multiple Instance Learning (MIL) offers a natural solution for settings where only coarse, bag-level labels are available, without having access to instance-level annotations. This is usually the case in digital pathology, which consists of gigapixel-sized images. While deterministic attention-based MIL approaches achieve strong bag-level performance, they often overlook the uncertainty inherent in instance relevance. In this paper, we address the lack of uncertainty quantification in instance-level attention scores by introducing SGPMIL, a new probabilistic attention-based MIL framework grounded in Sparse Gaussian Processes (SGP). By learning a posterior distribution over attention scores, SGPMIL enables principled uncertainty estimation, resulting in more reliable and calibrated instance relevance maps. Our approach not only preserves competitive bag-level performance but also significantly improves the quality and interpretability of instance-level predictions under uncertainty. SGPMIL extends prior work by introducing feature scaling in the SGP predictive mean function, leading to faster training, improved efficiency, and enhanced instance-level performance. Extensive experiments on multiple well-established digital pathology datasets highlight the effectiveness of our approach across both bag- and instance-level evaluations. Our code is available at https://github.com/mandlos/SGPMIL.
@@ 1 Introduction
Building on recent progress in MIL for WSI analysis, embedding-based MIL models with attention pooling have demonstrated strong performance at the bag-level while offering interpretable heatmaps that highlight task-relevant tissue regions [23, 43, 59, 56]. These methods rely on attention mechanisms to weigh instance contributions, yet the resulting scores are often interpreted heuristically as indicators of instance relevance. In this work, we introduce a probabilistic formulation of the attention mechanism by learning a posterior distribution over attention weights. This formulation enables uncertainty estimation and enhances the interpretability of instance-level predictions, while maintaining strong bag-level performance.
@@
Recent attention-based MIL aggregation schemes have demonstrated impressive bag-level performance across various metrics and datasets [23, 32, 43, 59, 12]. These results have been further enhanced by leveraging recent foundation models, which provide powerful instance-level representations. The aforementioned MIL works showcase attention heatmaps of test set cases for explainability purposes, where these heatmaps seem to focus on instances of interest which, for example, in a binary classification setting indicate the presence or absence of a positive class instance. While the attained bag-level performance is impressive, attention values are often interpreted as class-specific instance discriminators. We hypothesize that explicitly modeling uncertainty over these attention scores—by learning a posterior distribution—can improve instance-level performance, while preserving strong bag-level accuracy.
@@
Probabilistic deep learning offers principled tools for uncertainty estimation and improved model calibration [17], with methods such as Batch-Ensemble [54, 10], Concrete Dropout [15], and multiplicative Gaussian noise [26, 35, 8] approximating posterior distributions through ensembles, dropout, or noise injection. This is particularly valuable in MIL, where probabilistic frameworks can clarify prediction confidence, statistical properties, and the inner workings of pooling operations [56, 19, 12]. Prior works have explored uncertainty-driven MIL in various forms, including Bayesian CNNs under full supervision [47], Bayesian instance-based MIL under data scarcity and label noise [19, 36], uncertainty-driven aggregation models that refine attention or instance selection scores [56, 42, 52, 12, 46] and OOD uncertainty estimation [29]. Conformal prediction (CP) provides an alternative, distribution-free framework that yields uncertainty sets with marginal coverage guarantees, though it requires a separate calibration set and exchangeability assumptions [51]. Furthermore, Bayesian extensions of the embedding-based MIL have explored the performance/explainability gains that one has by inducing uncertainty to the attention scoring mechanism and imposing prior regularization [56, 42]. These approaches typically incorporate variational inference and Monte Carlo (MC) sampling to model distributional uncertainty at the instance and bag-level. Moreover, current challenges in probabilistic MIL approaches include numerical instability during training [42], limited evaluation on multiple datasets and tasks, or the need for an a priori decision on the range of instance spatial correlations in the scoring function [56].
@@
In this paper, we introduce SGPMIL, an efficient and robust probabilistic attention-based MIL framework grounded in Sparse Gaussian Processes (SGP). By designing a novel probabilistic formulation, our main contributions are structured along the following axes:
@@
(i) A fast and efficient probabilistic framework. SGPMIL builds upon prior work on SGP for MIL and introduces three key innovations: (1) a feature scaling term in the predictive mean of the learnable attention SGP variational posterior, which improves representational flexibility and adaptation to varying input scales; (2) a more stable and robust sampling strategy from the variational posterior during training, addressing numerical instabilities in Gaussian Process-based MIL approaches; and (3) an interpretable instance-level attention normalization mechanism across attention samples, enabling more consistent and meaningful instance-level predictions under the MIL setting.
@@
Extensive experiments across multiple datasets demonstrate on-par bag-level performance on binary classification tasks, with SGPMIL outperforming leading MIL approaches on multiclass datasets. Furthermore, we provide both intuitive motivation and empirical validation for the hypothesis that the inducing point approach in our MIL formulation captures salient histological morphologies and effectively focuses on task-relevant instances, leading to superior instance-level performance.
@@ 2.1 Deterministic frameworks
Many embedding-based MIL approaches compute instance-level attention scores through fixed parametric transformations. Ilse et al. [23] introduce attention-based deep MIL (ABMIL), where the slide-level representation is computed as a weighted average of instance embeddings. The attention weights are derived via a gated mechanism, allowing the model to learn instance relevance in a differentiable, end-to-end manner. Clustering-constrained attention MIL (CLAM) [32] extended the gated attention mechanism by introducing an extra clustering objective via a smooth SVM loss. Li et al. presented Dual-Stream MIL Network (DSMIL) [28], a dual-attention mechanism that combines instance and bag-level attention, where the most critical instance is selected via a classifier stream and its features guide the aggregation of the remaining patches. This is coupled with multiscale feature extraction and contrastive pretraining, leading to improved tumor detection performance. Transformer-based MIL (TransMIL) [43] made use of two self-attention heads [50] to weigh the instances and introduce spatial correlations using a novel Pyramid Position Encoding Generator (PPEG) achieving strong performance in bag-level classification metrics. Rymarczyk et al. [41] replace dot-product attention with kernel functions, underscoring the usefulness of kernelized self-attention for modeling instance correlations. More recent works like DGRMIL [59] model instance correlation by introducing cross-attention across instances and a set of learnable tokens which are trained by a positive instance alignment mechanism and a novel diversity loss term.
Despite strong bag-level performance, existing attention-based MIL methods rarely assess how well attention mechanisms identify task-relevant instances. Being deterministic, they also lack the capacity to estimate prediction uncertainty—crucial in medical applications. We address these gaps by introducing a probabilistic attention mechanism and conducting systematic instance-level evaluation, demonstrating improved identification of task-relevant instances along with principled uncertainty estimation.
@@ 2.2 Probabilistic frameworks
AGP [42] integrates a SGP layer within a multi-layer perceptron (MLP) to model attention distributions over instances for classification tasks. Given input embeddings, the SGP layer computes a variational multivariate normal posterior distribution over the attention values associated with each instance in a bag; Monte Carlo (MC) sampling yields multiple attention values, which are used to reweigh the instance embeddings, resulting in multiple samples of bag-level representations that are subsequently classified. The model is trained by maximizing the Evidence Lower Bound (ELBO) with respect to the model parameters. Another related approach is BayesMIL [56], which addresses the interpretability of attention mechanism in MIL by modeling probabilistic instance weights. Specifically, they introduce multiplicative Gaussian noise [26], into the projection layers and the classifier. To capture spatial correlations among instance attention distributions, they incorporate a conditional random field (CRF) on top of the learned attention distributions. MixMIL [12] integrates generalized linear mixed models with attention-based MIL by placing Gaussian priors on both instance and bag-level effects, injecting uncertainty into attention scores and estimating respective posterior distributions through variational inference; while robust and interpretable, its largely linear formulation may hinder performance on complex datasets.
Despite their promise, existing probabilistic approaches lack evaluations with pathology foundation encoders on cancer datasets, extensive instance-level evaluation on other datasets or rely on fixed spatial smoothing kernels, further requiring manual tuning [56]. AGP [42], in particular, faces numerical instabilities and scalability issues on larger datasets, and does not assess instance-level performance. In this work, we address these limitations by introducing a feature scaling term in the mean of the SGP variational posterior distribution, relaxed attention normalization and an efficient sampling strategy during training in AGP, addressing different instabilities, while conducting a comprehensive evaluation of both bag- and instance-level performance on the widely used public benchmarks.
@@
Gaussian Processes (GPs) offer a powerful, flexible, and elegant framework for regression, as is indicated by an extensive line of works [44, 25, 48]. As non-parametric standalone regressors, they mitigate the need for application-specific model architectures and a priori decision on model complexity [24, 55]. Furthermore, [38, 27] showcase the equivalence of GPs to an infinite-width single-layer fully-connected neural network with an i.i.d. prior over its parameters and with infinitely wide deep networks respectively which in turn are known to be universal function approximators [22]. Yet, these benefits come at the cost of time and storage complexity of the order of $\mathcal{O}(n^{3})$ and $\mathcal{O}(n^{2})$ respectively mainly due to inverting an $n\times n$ kernel matrix, $n$ being the number of training points, which is prohibitive for large datasets [37]. To overcome the time and storage cost bottleneck, Sparse Gaussian Processes were originally introduced by [44] and subsequently extended to allow for mini-batch training [21]. We leverage the scalable variational GPs framework introduced in [21] to induce uncertainty over instance attention distributions, and incorporate distributional correlation through kernel-based terms, showing that modeling such semantic correlations substantially improves instance-level performance.
@@ 3.1 Preliminaries
Figure 1: SGPMIL architecture overview. On the left, we illustrate the complete processing pipeline: WSIs are segmented and tiled into patches, which are then encoded using a frozen foundation model. The resulting patch embeddings are passed through an attention-based MIL head incorporating a probabilistic SGP mechanism. The attention-weighted embeddings are summed and projected through a trainable MLP. On the right, we highlight the probabilistic attention component. The SGP layer receives the embeddings along with learnable inducing points and infers a variational posterior over patch-level attention scores. Multiple attention samples are drawn to reweigh the embeddings, which are then aggregated into stochastic slide-level and classified via a linear layer followed by softmax.
Multiple Instance Learning (MIL) In MIL, the input is a bag of instances, $X_{i}=\{x_{1},\ldots,x_{K_{i}}\}$ with k-th instance $x_{k}\in\mathbb{R}^{D}$, associated with a bag label $Y_{i}\in\mathbb{N}$, while the instance labels $\{y_{1},\ldots,y_{K_{i}}\}$ remain unobserved during training; here $K_{i}$ denotes the number of instances in bag $i$. A dataset consists of $N$ bag-label pairs, denoted as $\mathcal{D}={(X_{i},Y_{i})}_{i=1}^{N}$, and let $\mathcal{C}=\{1,\ldots,C\}$ denote the set of class labels. The standard MIL assumption is defined as:
$Y_{i}=\begin{cases}0,&\text{if }\sum_{k=1}^{K_{i}}y_{k}=0\\[5.0pt] c,&\text{if }\exists\,k\text{ such that }y_{k}=c,c\in\mathcal{C}\end{cases}$ (1)
Sparse Gaussian Processes (SGPs) Following standard notation, given a training dataset $\{\left(X_{i},Y_{i}\right)\}_{i=1}^{n}$ of size $n$, where $Y_{i}=F_{i}+\epsilon_{i}$ is a noisy observation of the noise-free real-valued latent function $F(\cdot):X\in\mathcal{X}\rightarrow\mathbb{R}$ evaluated at $X_{i}$ and $\epsilon_{i}\sim\mathcal{N}\left(0,\sigma^{2}\right)$ models independent observation noise, an additional set of m learnable inducing points $\mathcal{Z}=\{Z_{i}\}_{i=1}^{m}$, $m\ll n$ are introduced to summarize the information of the entire dataset. These points $Z_{i}\in\mathcal{Z}$ reside in the same space as the inputs $X_{i}\in\mathcal{X}$ and are optimized during training to capture the underlying data distribution. In our formulation, $n$ is the total number of instances across all bags, $m$ is the number of learnable prototypical instance representations and we train the SGP layer using batches of instances, where each batch contains instances from the same bag. Let the vector $\mathbf{F}\in\mathbb{R}^{n}$ contain the values of function $F$ evaluated at each training point $X_{i}$ and the vector $\mathbf{U}\in\mathbb{R}^{m}$ be the values of the same function at the points $Z_{i}\in\mathcal{Z}$. Assuming that the vectors $(\mathbf{F},\mathbf{U})$ are jointly gaussian, the SGP model evaluated at finite sets of inputs $(\mathbf{X},\mathbf{Z})$ can be characterized by the following equations,
@@
[…]